# How Private is DP-SGD?

### v1

@@ Abstract
We demonstrate a substantial gap between the privacy guarantees of the Adaptive Batch Linear Queries (ABLQ) mechanism under different types of batch sampling: (i) Shuffling, and (ii) Poisson subsampling; the typical analysis of Differentially Private Stochastic Gradient Descent (DP-SGD) follows by interpreting it as a post-processing of ABLQ. While shuffling based DP-SGD is more commonly used in practical implementations, it is neither analytically nor numerically amenable to easy privacy analysis. On the other hand, Poisson subsampling based DP-SGD is challenging to scalably implement, but has a well-understood privacy analysis, with multiple open-source numerically tight privacy accountants available. This has led to a common practice of using shuffling based DP-SGD in practice, but using the privacy analysis for the corresponding Poisson subsampling version. Our result shows that there can be a substantial gap between the privacy analysis when using the two types of batch sampling, and thus advises caution in reporting privacy parameters for DP-SGD.
@@ 1 Introduction
Using noisy gradients in first order methods such as stochastic gradient descent has become a prominent approach for adding differential privacy ($\mathsf{DP}$) to the training of differentiable models such as neural networks. This approach, introduced by Abadi et al. (2016), has come to be known as Differentially Private Stochastic Gradient Descent, and we use the term DP-SGD to refer to any such first order method. DP-SGD is currently the canonical algorithm for training deep neural networks with privacy guarantees, and there currently exist multiple open source implementations, such as in Tensorflow Privacy (a), Pytorch Opacus (Yousefpour et al., 2021), and JAX Privacy (Balle et al., 2022). The algorithm has been applied widely in various machine learning domains, such as training of image classification (Tramer & Boneh, 2020; Papernot et al., 2021; Klause et al., 2022; De et al., 2022; Bu et al., 2022), generative models with GAN (Torkzadehmahani et al., 2019; Chen et al., 2020), diffusion models (Dockhorn et al., 2022), language models (Li et al., 2021; Yu et al., 2021; Anil et al., 2022; He et al., 2022), medical imaging (Ziller et al., 2021), as well as private spatial querying (Zeighami et al., 2021), ad modeling (Denison et al., 2023), and recommendation (Fang et al., 2022).
@@
Figure 1: Privacy parameter $\varepsilon$ for different noise parameters $\sigma$, for fixed $\delta=10^{-6}$ and number of steps $T=100,000$. $\varepsilon_{\mathcal{D}}$ : for deterministic batching, $\varepsilon_{\mathcal{P}}$ : upper bounds when using Poisson subsampling (computed using different accountants), and $\varepsilon_{\mathcal{S}}$ : a lower bound when using shuffling. We observe that shuffling does not provide much amplification for small values of $\sigma$, incurring significantly higher privacy cost compared to Poisson subsampling.
@@
DP-SGD operates by processing the training data in mini-batches, and at each step, performs a first order gradient update, using a noisy estimate of the average gradient for each mini-batch. In particular, the gradient $g$ for each record is first clipped to have a pre-determined bounded $\ell_{2}$-norm, by setting $[g]_{C}:=g\cdot\min\{1,C/\|g\|_{2}\}$, and then adding Gaussian noise of standard deviation $\sigma C$ to all coordinates of the sum of gradients in the mini-batch. [^1] While some distributed training setup uses the sum gradient directly, it is common to rescale the sum gradient by the batch size to obtain the average gradient before applying the optimization step. Since this scaling factor can be assimilated in the learning rate, we focus on the sum gradient for simplicity in this paper. The privacy guaranteed by the mechanism depends on the following: the choice of $\sigma$, the size of dataset, the number of steps of gradient update performed, and finally the process used to generate the batches. Almost all deep learning systems generate fixed-sized batches of data by going over the dataset sequentially. When feasible, a global shuffling of all the examples in the dataset is performed for each training epoch by making a single pass over the dataset. On the other hand, the process analyzed in Abadi et al. (2016) constructs each batch by including each record with a certain probability, chosen i.i.d. However, this leads to variable-sized mini-batches, which is technically challenging to handle in practice. As a result, there is generally a mismatch between the actual training pipeline and the privacy accounting in many applications of DP-SGD, with the implicit assumption that this subtle difference is negligible and excusable. However, in this paper, we show that this is not true—in a typical setting, as shown in Figure 1, the privacy loss from the correct accounting is significantly larger than expected.
@@
Formally, the privacy analysis of DP-SGD, especially in the case of non-convex differentiable models, is performed by viewing it as a post-processing of a mechanism performing adaptive batch linear queries ($\textsc{ABLQ}_{\mathcal{B}}$) as defined in Algorithm 1 (we use subscript $\mathcal{B}$ to emphasize the role of the batch sampler), where the linear query $\psi_{t}(x_{i})$ corresponds to the clipped gradient corresponding to record $x_{i}$. For any $x$, we require that $\psi_{t}(x)\in\mathbb{B}^{d}:=\{w\in\mathbb{R}^{d}:\|w\|_{2}\leq 1\}$. Note that without loss of generality, we can treat the norm bound $C=1$ by defining $\psi_{t}(x):=[g]_{C}/C$ for the corresponding gradient $g$, and rescaling $g_{t}$ by $C$ to get back the noisy gradient for DP-SGD.
As mentioned above, a canonical way to generate the mini-batches is to go through the dataset in a fixed deterministic order, and divide the data into mini-batches of a fixed size (Algorithm 2, denoted as $\mathcal{D}$). Another commonly used option is to first randomly permute the entire dataset before dividing it into mini-batches of a fixed size (Algorithm 3, denoted as $\mathcal{S}$); this option provides amplification by shuffling, namely, that the privacy guarantees are better compared to fixed deterministic ordering. However, obtaining such amplification bounds is non-trivial, and while some bounds have been recently established for such privacy amplification (Erlingsson et al., 2019; Feldman et al., 2021, 2023), they tend to be loose and only kick in when the basic mechanism is already sufficiently private.
@@
Instead, the approach followed by Abadi et al. (2016), and henceforth used commonly in reporting privacy parameters for DP-SGD, is to assume that each batch is sampled i.i.d. by including each record with a certain probability, referred to as Poisson subsampling (Algorithm 4, denoted as $\mathcal{P}$). The advantage of this approach is that the privacy analysis of such sampling is easy to carry out since the $\textsc{ABLQ}_{\mathcal{P}}$ mechanism can be viewed as a composition of $T$ independent sub-mechanisms. This enables privacy accounting methods such as Rényi DP (Mironov, 2017) as well as numerically tight accounting methods using privacy loss distributions (elaborated on later in Section 3).
This has been the case, even when the algorithm being implemented in fact uses some form of shuffling based batch sampler. To quote Abadi et al. (2016) (with our emphasis),
@@
Most implementations of DP-SGD mentioned earlier also use some form of shuffling, with a rare exception of PyTorch Opacus (Yousefpour et al., 2021) that has the option of Poisson subsampling to be consistent with the privacy analysis. But this approach does not scale to large datasets as random access for datasets that do not fit in memory is generally inefficient. Moreover, variable batch sizes are inconvenient to handle in deep learning systems [^2] For example, when the input shape changes, jax.jit will trigger recompilation, and tf.function will retrace the graph. Google TPUs require all operations to have fixed (input and output) shapes. Moreover, in various form of data parallelism, the batch size needs to be divisible by the number of accelerators.. Tensorflow Privacy (b) provides the compute_dp_sgd_privacy_statement method for computing the privacy parameters and reminds users about the implicit assumption of Poisson subsampling. Ponomareva et al. (2023) note in their survey that “It is common, though inaccurate, to train without Poisson subsampling, but to report the stronger DP bounds as if amplification was used.”
As DP-SGD is being deployed in more applications with such discrepancy, it has become crucial to understand exactly how the privacy guarantee depends on the precise choice of the batch sampler, especially when one cares about specific $(\varepsilon,\delta)$ privacy parameters (and not asymptotic bounds). This leads us to the main question motivating this work:
@@ 1.1 Our Contributions
We show that the privacy guarantee of $\textsc{ABLQ}_{\mathcal{D}}$ and $\textsc{ABLQ}_{\mathcal{P}}$ are incomparable. Namely, for all values of $T$ (number of steps) and $\sigma$ (noise parameter), it holds (i) for small enough $\varepsilon>0$ that $\delta_{\mathcal{P}}(\varepsilon)<\delta_{\mathcal{D}}(\varepsilon)$, but perhaps more surprisingly, (ii) for sufficiently large $\varepsilon$ it holds that $\delta_{\mathcal{P}}(\varepsilon)\gg\delta_{\mathcal{D}}(\varepsilon)$. We also demonstrate this separation in a specific numerical setting of parameters.
@@
For example, in Figure 1, we fix $\delta=10^{-6}$ and the number of steps $T=100,000$, and compare the value of $\varepsilon_{\mathcal{B}}(\delta)$ for various values of $\sigma$. In particular, for $\sigma=0.4$, we find that $\varepsilon_{\mathcal{P}}(\delta)\leq 3$ which could be considered as a reasonable privacy guarantee, but $\varepsilon_{\mathcal{S}}(\delta)\geq 14.45$.
This suggests that reporting privacy guarantees using the Poisson batch sampler can significantly underestimate the privacy loss when in fact the implementation uses the shuffle batch sampler.
Our main takeaway is that batch sampling plays a crucial in determining the privacy guarantees of $\textsc{ABLQ}_{\mathcal{B}}$, and hence caution must be exercised in reporting privacy parameters for mechanisms such as DP-SGD.
@@ 1.2 Technical Overview
We demonstrate the separation in a specific numerical settings of parameters by using the dp_accounting library (Google’s DP Library., 2020) to provide lower bounds on the hockey stick divergence between the tightly dominating pair for $\textsc{ABLQ}_{\mathcal{P}}$.
@@
We provide the IPython notebook (Google Colab link) that was used for all the numerical demonstrations; the entire notebook executes in a few minutes when using a free CPU runtime on Google Colab .
@@ 2 Differential Privacy
We consider mechanisms that map input datasets to distributions over an output space, namely $\mathcal{M}:\mathcal{X}^{*}\to\Delta_{\mathcal{O}}$. That is, on input dataset $\bm{x}=(x_{1},\ldots,x_{n})$ where each record $x_{i}\in\mathcal{X}$, $\mathcal{M}(\bm{x})$ is a probability distribution over the output space $\mathcal{O}$. Two datasets $\bm{x}$ and $\bm{x}^{\prime}$ are said to be adjacent, denoted $\bm{x}\sim\bm{x}^{\prime}$, if, loosely speaking, they “differ in one record”. This can be formalized in multiple ways, which we elaborate on in Section 2.2, but for any notion of adjacency, Differential Privacy (DP) can be defined as follows.
@@ 2.1 Adaptive Batch Linear Queries Mechanism
We primarily study the adaptive batch linear queries mechanism $\textsc{ABLQ}_{\mathcal{B}}$ using a batch sampler $\mathcal{B}$ and an adaptive query method $\mathcal{A}$, as defined in Algorithm 1. The batch sampler $\mathcal{B}$ can be instantiated with any algorithm that produces a (randomized) sequence of batches $S_{1},\ldots,S_{T}$. $\textsc{ABLQ}_{\mathcal{B}}$ produces a sequence of responses $(g_{1},\ldots,g_{T})$ where each $g_{i}\in\mathbb{R}^{d}$. The response $g_{t}$ is produced recursively using the adaptive query method $\mathcal{A}$ that given $g_{1},\ldots,g_{t-1}$, constructs a new query $\psi_{t}:\mathcal{X}\to\mathbb{B}^{d}$ (for $\mathbb{B}^{d}:=\{v\in\mathbb{R}^{d}:\|v\|_{2}\leq 1\}$), and we estimate the average value of $\psi_{t}(x)$ over the batch $S_{t}$ with added zero-mean Gaussian noise of scale $\sigma$ to all coordinates. As explained in Section 1, DP-SGD falls under this abstraction by considering an adaptive query method that is specified by a differentiable loss function $f:\mathbb{R}^{d}\times\mathcal{X}\to\mathbb{R}$, and starts with an initial $w_{0}\in\mathbb{R}^{d}$, and defines the query $\mathcal{A}(g_{1},\ldots,g_{t-1};x)$ as the clipped gradient $[\nabla_{w}f_{w_{t-1}}(x)]_{1}$ where $w_{t}$ is the $t$-th model iterate recursively obtained by performing gradient descent, e.g., $w_{t}\leftarrow w_{t-1}-\eta_{t}g_{t}$ (or any other first order optimization step).

### v2

# How Private are DP-SGD Implementations?

@@ Abstract
We demonstrate a substantial gap between the privacy guarantees of the Adaptive Batch Linear Queries (ABLQ) mechanism under different types of batch sampling: (i) Shuffling, and (ii) Poisson subsampling; the typical analysis of Differentially Private Stochastic Gradient Descent (DP-SGD) follows by interpreting it as a post-processing of ABLQ. While shuffling-based DP-SGD is more commonly used in practical implementations, it has not been amenable to easy privacy analysis, either analytically or even numerically. On the other hand, Poisson subsampling-based DP-SGD is challenging to scalably implement, but has a well-understood privacy analysis, with multiple open-source numerically tight privacy accountants available. This has led to a common practice of using shuffling-based DP-SGD in practice, but using the privacy analysis for the corresponding Poisson subsampling version. Our result shows that there can be a substantial gap between the privacy analysis when using the two types of batch sampling, and thus advises caution in reporting privacy parameters for DP-SGD.
@@ 1 Introduction
Using noisy gradients in first-order methods such as stochastic gradient descent (SGD) has become a prominent approach for adding differential privacy ($\mathsf{DP}$) to the training of differentiable models such as neural networks. This approach, introduced by Abadi et al. (2016), has come to be known as Differentially Private Stochastic Gradient Descent, and we use the term $\mathsf{DP}\text{-}\mathsf{SGD}$ to refer to any such first-order method. $\mathsf{DP}\text{-}\mathsf{SGD}$ is currently the canonical algorithm for training deep neural networks with privacy guarantees, and there currently exist multiple open source implementations, such as in Tensorflow Privacy (a), Pytorch Opacus (Yousefpour et al., 2021), and JAX Privacy (Balle et al., 2022). The algorithm has been applied widely in various machine learning domains, such as training of image classification (Tramer & Boneh, 2021; Papernot et al., 2021; Klause et al., 2022; De et al., 2022; Bu et al., 2022), generative models with GAN (Torkzadehmahani et al., 2019; Chen et al., 2020), diffusion models (Dockhorn et al., 2022), language models (Li et al., 2022; Yu et al., 2022; Anil et al., 2022; He et al., 2023), medical imaging (Ziller et al., 2021), as well as private spatial querying (Zeighami et al., 2022), ad modeling (Denison et al., 2023), and recommendation (Fang et al., 2022).
@@
Figure 1: Privacy parameter $\varepsilon$ for different noise parameters $\sigma$, for fixed $\delta=10^{-6}$ and number of steps $T=10,000$. $\varepsilon_{\mathcal{D}}$ : for deterministic batching, $\varepsilon_{\mathcal{P}}$ : upper bounds when using Poisson subsampling (computed using different accountants), and $\varepsilon_{\mathcal{S}}$ : a lower bound when using shuffling. We observe that shuffling does not provide much amplification for small values of $\sigma$, incurring significantly higher privacy cost compared to Poisson subsampling.
@@
$\mathsf{DP}\text{-}\mathsf{SGD}$ operates by processing the training data in mini-batches, and at each step, performs a first-order gradient update, using a noisy estimate of the average gradient for each mini-batch. In particular, the gradient $g$ for each record is first clipped to have a pre-determined bounded $\ell_{2}$-norm, by setting $[g]_{C}:=g\cdot\min\{1,C/\|g\|_{2}\}$, and then adding Gaussian noise of standard deviation $\sigma C$ to all coordinates of the sum of gradients in the mini-batch. [^1] While some distributed training setup uses the sum gradient directly, it is common to rescale the sum gradient by the batch size to obtain the average gradient before applying the optimization step. Since this scaling factor can be assimilated in the learning rate, we focus on the sum gradient for simplicity in this paper. Note that, in case of $\mathsf{DP}\text{-}\mathsf{SGD}$ using Poisson subsampling, the scaling is done by the expected batch size, and not the realized batch size. The privacy guaranteed by the mechanism depends on the following: the choice of $\sigma$, the size of dataset, the size of mini-batches, the number of steps of gradient update performed, and finally the process used to generate the batches. Almost all deep learning systems generate fixed-sized batches of data by going over the dataset sequentially. When feasible, a global shuffling of all the examples in the dataset is performed for each training epoch by making a single pass over the dataset. On the other hand, the process analyzed by Abadi et al. (2016) constructs each batch by including each record with a certain probability, chosen i.i.d. However, this leads to variable-sized mini-batches, which is technically challenging to handle in practice. As a result, there is generally a mismatch between the actual training pipeline and the privacy accounting in many applications of $\mathsf{DP}\text{-}\mathsf{SGD}$, with the implicit assumption that this subtle difference is negligible and excusable. However, in this paper, we show that this is not true—in a typical setting, as shown in Figure 1, the privacy loss from the correct accounting is significantly larger than expected.
@@
Formally, the privacy analysis of $\mathsf{DP}\text{-}\mathsf{SGD}$, especially in the case of non-convex differentiable models, is performed by viewing it as a post-processing of a mechanism performing adaptive batch linear queries ($\textsc{ABLQ}_{\mathcal{B}}$) as defined in Algorithm 1 (we use subscript $\mathcal{B}$ to emphasize the role of the batch sampler), where the linear query $\psi_{t}(x_{i})$ corresponds to the clipped gradient corresponding to record $x_{i}$. For any $x$, we require that $\psi_{t}(x)\in\mathbb{B}^{d}:=\{w\in\mathbb{R}^{d}:\|w\|_{2}\leq 1\}$. Note that without loss of generality, we can treat the norm bound $C$ to be $1$ by defining $\psi_{t}(x):=[g]_{C}/C$ for the corresponding gradient $g$, and rescaling $g_{t}$ by $C$ to get back the noisy gradient for $\mathsf{DP}\text{-}\mathsf{SGD}$.
As mentioned above, a canonical way to generate the mini-batches is to go through the dataset in a fixed deterministic order, and divide the data into mini-batches of a fixed size (Algorithm 2, denoted as $\mathcal{D}$). Another commonly used option is to first randomly permute the entire dataset before dividing it into mini-batches of a fixed size (Algorithm 3, denoted as $\mathcal{S}$); this option provides amplification by shuffling, namely, that the privacy guarantees are better compared to fixed deterministic ordering. However, obtaining such amplification bounds is non-trivial, and while some bounds have been recently established for such privacy amplification (Erlingsson et al., 2019; Feldman et al., 2021, 2023), they tend to be loose in our setting and only kick in when the basic mechanism is already sufficiently private.
@@
Instead, the approach followed by (Abadi et al., 2016), and henceforth used commonly in reporting privacy parameters for $\mathsf{DP}\text{-}\mathsf{SGD}$, is to assume that each batch is sampled i.i.d. by including each record with a certain probability, referred to as Poisson subsampling (Algorithm 4, denoted as $\mathcal{P}$). The advantage of this approach is that the privacy analysis of such sampling is easier to carry out since the $\textsc{ABLQ}_{\mathcal{P}}$ mechanism can be viewed as a composition of $T$ independent sub-mechanisms. This enables privacy accounting methods such as Rényi DP (Mironov, 2017) as well as numerically tight accounting methods using privacy loss distributions (elaborated on later in Section 3).
This has been the case, even when the algorithm being implemented in fact uses some form of shuffling-based batch sampler. To quote Abadi et al. (2016) (with our emphasis),
@@
Most implementations of $\mathsf{DP}\text{-}\mathsf{SGD}$ mentioned earlier also use some form of shuffling, with a rare exception of PyTorch Opacus (Yousefpour et al., 2021) that has the option of Poisson subsampling to be consistent with the privacy analysis. But this approach does not scale to large datasets as random access for datasets that do not fit in memory is generally inefficient. Moreover, variable batch sizes are inconvenient to handle in deep learning systems [^2] For example, when the input shape changes, jax.jit will trigger recompilation, and tf.function will retrace the graph. Google TPUs require all operations to have fixed (input and output) shapes. Moreover, in various form of data parallelism, the batch size needs to be divisible by the number of accelerators.. Tensorflow Privacy (b) provides the compute_dp_sgd_privacy_statement method for computing the privacy parameters and reminds users about the implicit assumption of Poisson subsampling. Ponomareva et al. (2023) note in their survey that “It is common, though inaccurate, to train without Poisson subsampling, but to report the stronger DP bounds as if amplification was used.”
As $\mathsf{DP}\text{-}\mathsf{SGD}$ is being deployed in more applications with such discrepancy, it has become crucial to understand exactly how the privacy guarantee depends on the precise choice of the batch sampler, especially when one cares about specific $(\varepsilon,\delta)$ privacy parameters (and not asymptotic bounds). This leads us to our main motivating question:
@@ 1.1 Our Contributions
We show that the privacy guarantee of $\textsc{ABLQ}_{\mathcal{D}}$ and $\textsc{ABLQ}_{\mathcal{P}}$ are incomparable. Namely, for all values of $T$ (number of steps) and $\sigma$ (noise parameter), it holds (i) for small enough $\varepsilon>0$ that $\delta_{\mathcal{P}}(\varepsilon)<\delta_{\mathcal{D}}(\varepsilon)$, but perhaps more interestingly, (ii) for sufficiently large $\varepsilon$ it holds that $\delta_{\mathcal{P}}(\varepsilon)\gg\delta_{\mathcal{D}}(\varepsilon)$. We also demonstrate this separation in a specific numerical setting of parameters.
@@
For example, in Figure 1, we fix $\delta=10^{-6}$ and the number of steps $T=10,000$, [^3] Recall that since we are analyzing a “single epoch”, the subsampling probability of Poisson subsampling is $b/n=1/T$. and compare the value of $\varepsilon_{\mathcal{B}}(\delta)$ for various values of $\sigma$. For $\sigma=0.5$, we find $\varepsilon_{\mathcal{P}}(\delta)<1.96$ (PLD) and $\varepsilon_{\mathcal{P}}(\delta)<3.43$ (RDP), but $\varepsilon_{\mathcal{S}}(\delta)>10.994$ and $\varepsilon_{\mathcal{D}}(\delta)\approx 10.997$. For $\sigma=1.3$, we find $\varepsilon_{\mathcal{P}}(\delta)<0.031$ (PLD), whereas, $\varepsilon_{\mathcal{S}}(\delta)>0.26$.
This suggests that reporting privacy guarantees using the Poisson batch sampler can significantly underestimate the privacy loss when the implementation in fact uses the shuffle batch sampler.
Our main takeaway is that batch sampling plays a crucial in determining the privacy guarantees of $\textsc{ABLQ}_{\mathcal{B}}$, and hence caution must be exercised in reporting privacy parameters for mechanisms such as $\mathsf{DP}\text{-}\mathsf{SGD}$.
@@ 1.2 Technical Overview
We demonstrate the separation in a specific numerical setting of parameters using the dp_accounting library (Google’s DP Library., 2020) to provide lower bounds on the hockey stick divergence between the tightly dominating pair for $\textsc{ABLQ}_{\mathcal{P}}$.
@@
We provide the IPython notebook [^4] https://colab.research.google.com/drive/1zI2H8YEXbQyD6gZVVskFwcOiM5YMvqRe?usp=sharing that was used for all the numerical demonstrations; the notebook can be executed using a free CPU runtime on Google Colab .
Related work.
The phenomenon of non-differing records leaking information about whether the differing record is in a batch or not can also exist in sampling-based batch samplers, such as, sampling independent batches of fixed-size, as studied in a concurrent work by Lebeda et al. (2024). However, we focus on the shuffle-based batch sampler, as these are most common in practical implementations.
@@ 2 Differential Privacy
We consider mechanisms that map input datasets to distributions over an output space, namely $\mathcal{M}:\mathcal{X}^{*}\to\Delta_{\mathcal{O}}$. That is, on input dataset $\bm{x}=(x_{1},\ldots,x_{n})$ where each record $x_{i}\in\mathcal{X}$, $\mathcal{M}(\bm{x})$ is a probability distribution over the output space $\mathcal{O}$; we abuse notation to also use $\mathcal{M}(\bm{x})$ to denote the underlying random variable. Two datasets $\bm{x}$ and $\bm{x}^{\prime}$ are said to be adjacent, denoted $\bm{x}\sim\bm{x}^{\prime}$, if, loosely speaking, they “differ in one record”. This can be formalized in multiple ways, which we elaborate on in Section 2.2, but for any notion of adjacency, Differential Privacy (DP) can be defined as follows.
@@ 2.1 Adaptive Batch Linear Queries Mechanism
We primarily study the adaptive batch linear queries mechanism $\textsc{ABLQ}_{\mathcal{B}}$ using a batch sampler $\mathcal{B}$ and an adaptive query method $\mathcal{A}$, as defined in Algorithm 1. The batch sampler $\mathcal{B}$ can be instantiated with any algorithm that produces a (randomized) sequence $S_{1},\ldots,S_{T}\subseteq[n]$ of batches, where $n$ is the number of examples. $\textsc{ABLQ}_{\mathcal{B}}$ produces a sequence $(g_{1},\ldots,g_{T})$ of responses where each $g_{i}\in\mathbb{R}^{d}$. The response $g_{t}$ is produced recursively using the adaptive query method $\mathcal{A}$ that given $g_{1},\ldots,g_{t-1}$, constructs a new query $\psi_{t}:\mathcal{X}\to\mathbb{B}^{d}$ (for $\mathbb{B}^{d}:=\{v\in\mathbb{R}^{d}:\|v\|_{2}\leq 1\}$), and we estimate the sum of $\psi_{t}(x)$ over the batch $S_{t}$ with added zero-mean Gaussian noise of scale $\sigma$ to all coordinates. As explained in Section 1, $\mathsf{DP}\text{-}\mathsf{SGD}$ falls under this abstraction by considering an adaptive query method that is specified by a differentiable loss function $f:\mathbb{R}^{d}\times\mathcal{X}\to\mathbb{R}$, and starts with an initial $w_{0}\in\mathbb{R}^{d}$, and defines the query $\mathcal{A}(g_{1},\ldots,g_{t-1};x)$ as the clipped gradient $[\nabla_{w}f_{w_{t-1}}(x)]_{1}$ where $w_{t}$ is the $t$th model iterate recursively obtained by performing gradient descent, e.g., $w_{t}\leftarrow w_{t-1}-\eta_{t}g_{t}$ (or any other first-order optimization step).